OpenStudy (anonymous):

x = 2y - 3 I need to find x intercept and y intercept. I have (-3,0)(0,3/2) Is this correct?

OpenStudy (anonymous):

on my homework sheet the equation is: x = 2y - 3

OpenStudy (anonymous):

I thought to figure the x intercept you change the y to a 0 and then to find the y intercept you change the x to a 0. Is that wrong? I didn't know I needed thee slope intercept formula. And no I don't believe I know the slope intercept formula.. I would have to look it up

OpenStudy (anonymous):

Yes, please look up and memorize it if it's possible to you!

OpenStudy (anonymous):

do you mean that (-3,0)(0,3/2) is not correct then?

OpenStudy (anonymous):

ok the slope intercept formula is: y = mx + b

OpenStudy (anonymous):

how do I apply the equation on my homework sheet to that?? x = 2y - 3

OpenStudy (anonymous):

Yep, now arrange your equation into this order: y = mx + b

OpenStudy (anonymous):

-2y = 3 - x is that how it should be set up?

OpenStudy (anonymous):

I mean y = x/2 + 3/2

OpenStudy (anonymous):

Oops, I understand what you means now, there's no need to change into point slope form to find x intercept and y intercept !

OpenStudy (anonymous):

I don't understand... I will have to think abuot this

OpenStudy (anonymous):

whew!!! so we were headed down different paths????

OpenStudy (anonymous):

Your explanation about finding x, y intercept is absolutely right!

OpenStudy (anonymous):

Thank you !! I truly appreciate your help!!

OpenStudy (anonymous):

more than one way to solve a problem!!!!

OpenStudy (anonymous):

You can see clearly in the Slope Intercept form that y intercept = 3/2 ( 0, 3/2)

OpenStudy (anonymous):

y = mx + b b is y intercept 3/2

OpenStudy (anonymous):

yes - I believe I understand what you are saying

OpenStudy (anonymous):

Your result is correct at the first place :)

OpenStudy (anonymous):

I do appreciate your help

OpenStudy (anonymous):

Still the Slope Intercept form is extremely essential for algebra learning, so memorize and know how to apply it!

OpenStudy (anonymous):

will do!!

OpenStudy (anonymous):

You probably will learn it in 2 weeks at most!